OpenStudy (anonymous):

Ok this is what I have h=(4,38) &p=(13,95) I need to fine the point slope of an equation of the line through the 2 ordered pairs. My options are: a.) p=19/3h+38/3 b.) p-4=3/19(h-38) c.) p-38=19/3(h-4) d.) p-38=-19/3(h-4) so which one is the answer?

OpenStudy (anonymous):

first find the slope by using this formula (y2-yz)/(x2-x) or (95-38)/(13-4) = 57/9=19/3 so it can't be b or d

OpenStudy (anonymous):

then substitue into the slope-point form: y - y1=m(x-x1) y-38=19/3(x-4) so it has to be c

OpenStudy (anonymous):

thanks so no it is asking to use that formula to estimate the production of an 18 foot tree anda 20 foot tree